OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 4x - 2y if x = -1 and y = -4

OpenStudy (anonymous):

4

OpenStudy (anonymous):

4(-1) - 2(-4) = -4 + 8 = 4